# Attention Surgery: An Efficient Recipe to Linearize Your Video Diffusion Transformer

### v2

@@ Abstract
Transformer-based video diffusion models (VDMs) deliver state-of-the-art video generation quality but are constrained by the quadratic cost of self-attention, making long sequences and high resolutions computationally expensive. While linear attention offers sub-quadratic complexity, prior attempts fail to match the expressiveness of softmax attention without costly retraining. We introduce Attention Surgery, an efficient framework for linearizing or hybridizing attention in pretrained VDMs without training from scratch. Inspired by recent advances in language models, our method combines a novel hybrid attention mechanism—mixing softmax and linear tokens—with a lightweight distillation and fine-tuning pipeline requiring only a few GPU-days. Additionally, we incorporate a cost-aware block-rate strategy to balance expressiveness and efficiency across layers. Applied to Wan2.1 1.3B, a state-of-the-art DiT-based VDM, Attention Surgery achieves the first competitive sub-quadratic attention video diffusion models, reducing attention cost by up to 40% in terms of FLOPs, while maintaining generation quality as measured on the standard VBench and VBench-2.0 benchmarks. Project page is available at: https://qualcomm-ai-research.github.io/attention-surgery
@@ 1 Introduction
Video diffusion models (VDMs) have become a cornerstone of generative modeling, enabling high-fidelity, temporally coherent video synthesis for applications from entertainment to simulation. Early VDMs relied on U-Net backbones, but these architectures struggle to scale and capture long-range temporal dependencies. Recent advances favor Diffusion Transformers (DiTs) (Peebles & Xie, 2023). These models operate on spatiotemporal patches and provide global receptive fields from the outset. State-of-the-art systems such as Wan2.1 (Wan et al., 2025), CogVideoX (Yang et al., 2025), HunyuanVideo (Lab, 2025), PyramidalFlow (Jin et al., 2025), and Open-Sora Plan (Lin et al., 2024) exemplify this trend, consistently outperforming U-Net-based models in quality and scalability. Recent surveys confirm this transition, highlighting DiTs as the dominant architecture for video generation (Wang et al., 2025b; Melnik et al., 2024).
While recent DiT-based video diffusion models deliver state-of-the-art quality, they come with substantial computational and memory costs that limit their practical applicability. A primary bottleneck lies in the self-attention mechanism, whose complexity scales quadratically with the sequence length, i.e., in time and in memory, where denotes the number of tokens and the hidden dimension (Vaswani et al., 2017; Rabe & Staats, 2021). This issue is particularly severe in video diffusion, where the token count easily reaches tens of thousands due to the combination of spatial patches and multiple frames. Our profiling of large-scale DiT-based video diffusion models indicates that a vast proportion of compute is devoted to self-attention. For instance, in Wan2.1 1.3B, more than 76% of the total compute within the transformer blocks is attributable to self-attention alone. Even with optimizations such as FlashAttention (Dao et al., 2022), the quadratic scaling remains a fundamental barrier, constraining both training and inference when targeting higher resolutions, longer durations, or multi-shot videos. Consequently, reducing the attention cost without sacrificing model quality is critical for making video diffusion models more efficient and broadly deployable.
@@
Although linear-time attention to address the quadratic cost of softmax attention (Katharopoulos et al., 2020; Choromanski et al., 2021) has been around for several years, no competitive video diffusion model based on linear attention exists today. Three factors explain this gap. First, training such models from scratch is prohibitively expensive: state-of-the-art video diffusion systems require hundreds of thousands to millions of GPU hours and massive curated datasets (Blattmann et al., 2023; Chen et al., 2024; Wan et al., 2025). Second, there is no practical method proposed for distilling softmax attention into linear attention under reasonable compute budgets for video diffusion. This difficulty arises because the exponential kernel underlying softmax requires an infinite-dimensional feature map for exact representation, making efficient approximations challenging (Han et al., 2024; Zhang et al., 2025). While linear attention in image diffusion (Li et al., 2023) and low-rank linearization in language models (Zhang et al., 2025) show promise, no analogous solution exists for the more complicated spatiotemporal token interactions in video diffusion. Third, lower expressiveness in linear attention often results in notable degradation in the attention transformation fidelity and consequently lower quality generations, specifically for videos with more complicated temporal signal dynamics.
To address these challenges, we propose an efficient attention surgery strategy – eliminating the need for extensive retraining from scratch – coupled with a novel efficient hybrid attention architecture inspired by recent developments in language modeling (Zhang et al., 2025). Intuitively, if a small subset of tokens retains full softmax attention while the rest use linear attention, the model can preserve global structure and fine-grained dependencies where needed, while scaling efficiently elsewhere. Our approach significantly narrows the quality gap between linearized and full softmax attention while achieving higher efficiency than the original softmax attention models. Importantly, it can be realized with modest compute –requiring less than 0.4k GPU hours for the overall surgery – making it practical for a wide range of research and industrial settings. We validate our method on Wan2.1, a state-of-the-art video diffusion model, demonstrating that our contributions are successfully applicable to transformer-based diffusion models. Figure 1 illustrates the obtained advantage.
@@ 2 Related Work
Efficient Attention. Numerous approaches have been proposed to reduce the quadratic complexity of self-attention, for perception e.g. EfficientViT (Cai et al., 2023), PADRe (Letourneau et al., 2025), Performer (Choromanski et al., 2021), and Linformer (Wang et al., 2020), for image generation e.g. SANA (Li et al., 2023), LinGen Wang et al. (2025a), Grafting (Chandrasegaran et al., 2025), and for language modeling (Mercat et al., 2024; Wang et al., 2024; Yang et al., 2024; Zhang et al., 2024; 2025). Although these methods demonstrate the feasibility of sub-quadratic or more efficient attention, they typically require extensive training or training from scratch (e.g., SANA (Li et al., 2023)). In contrast, our work aims for lightweight distillation and fine-tuning of pre-trained softmax-attention-based models into an efficient hybrid attention design specifically tailored for video diffusion under modest compute budgets. Moreover, we target video generation task with unique challenges in modeling spatiotemporal token dependencies.
@@
Linear recurrent models, such as SSM and RWKV, have recently emerged as efficient alternatives to self-attention, enabling the modeling of longer token sequences for high-resolution image generation (Fei et al., 2024b; Wang et al., 2024; Fei et al., 2024a; Zhu et al., 2025; Yao et al., 2025). However, due to the architectural differences between transformer blocks and SSM-based blocks (e.g., Mamba), distilling pre-trained DiT weights into these architectures typically requires extensive training. In contrast, our approach preserves the same underlying block structure, enabling effective distillation under a low-training regime. Furthermore, as highlighted in the seminal work on linear attention (Katharopoulos et al., 2020), there exists a strong connection between RNNs and causal linear attention, which allows a linearized causal attention mechanism to be deployed as an RNN during inference, a property that is particularly desirable for long video generation.
@@
Efficient Video Diffusion Models. Recent large-scale video diffusion systems such as CogVideoX (Yang et al., 2025), Open-Sora Plan (Lin et al., 2024), PyramidalFlow (Jin et al., 2025), LTX-video (HaCohen et al., 2024), and Wan2.1 (Wan et al., 2025) have advanced quality and scalability, but at enormous compute and memory cost. Mobile-oriented designs like Mobile Video Diffusion (Yahia et al., 2025), MoViE (Karjauv et al., 2024), SnapGen-V (Wu et al., 2025b; a), AMD-HummingBird (Isobe et al., 2025), and On-device Sora (Kim et al., 2025) explore lightweight architectures, yet remain non-DiT-based or rely on full quadratic attention, limiting scalability to long videos.
Prior work accelerates video generation via token merging (Bolya & Hoffman, 2023; Kahatapitiya et al., 2024; Ding et al., 2025), token downsampling (Crowson et al., 2024; Peruzzo et al., 2025), and attention tiling (Ding et al., 2025). Efficient-vDiT (Ding et al., 2025) improves DiT efficiency by tiling attention, achieving linear complexity and reducing memory overhead. In contrast, our hybrid attention with attention surgery remains quadratic in the worst case but reduces effective compute by applying softmax attention to only a fraction of tokens and linear attention to the rest. While not matching the asymptotic complexity of Efficient-vDiT, our method achieves better quality–efficiency trade-offs. Concurrent work M4V (Huang et al., 2025) accelerates video DiTs by distilling them into Mamba blocks. Despite our simpler block structure and lightweight training, we outperform M4V in both generation quality and efficiency.
@@ 3.1 Preliminaries: Linear Attention
As observed, the terms and are independent of the output index and thus can be precomputed and cached to achieve the above-defined reformulated attention in linear complexity. Note that must be non-negative and the original linear attention paper by Katharopoulos et al. (2020) defines it as , however, the significant mismatch in expressiveness of the original similarity function and the elu-based results in substantial retraining compute and data requirement and/or inability to achieve the original quality observed from softmax attention.
@@ 3.2 Hybrid Attention
Attention Architecture. To circumvent the aforementioned issues, and inspired by the recent developments in language models (Zhang et al., 2025), we define the hybrid attention by decoupling the full set of token indices into softmax tokens and the rest as linear tokens, , as:
Here is the stabilizing constant computed as the maximum exponent. As opposed to Zhang et al. (2025), instead of defining as a local window around token , we uniformly subsample tokens at hybrid rate , as: . This design ensures that higher-quality softmax tokens are distributed across the entire temporal span, providing global anchors that preserve motion coherence and prevent temporal drift — an issue that local windows often suffer from in video generation. Note that selecting subsampling rates is an important hyperparameter that indicates a trade-off between the fidelity of hybrid reconstruction of the original softmax and the corresponding computational efficiency. Higher values such as or will ensure that the attention to only a small fraction of tokens scales quadratically that will notably decrease the compute burden. In contrast, a value of can reconstruct the original softmax with higher accuracy, while still spending the quadratic terms on half of the tokens. Fig. 2 illustrates this.
@@ 3.3 Attention Surgery
where the logarithmic term mitigates numerical instabilities caused by large attention logits and gradients (Barron, 2025). However, given that matching the attention scores is a proxy optimization and the self-attention’s weighted averaged hidden states are the target to match, one can alternatively define the value distillation objective as follows:
@@
Heterogeneous Block-Rate Optimization. Referring to Fig. 3, we observe that various blocks exhibit different reconstruction error values under different hybrid rates, each with their corresponding compute costs. How do we then decide upon the hybrid rates (8, 2, 4 or 1 (full softmax)) to form our hybrid architecture among the exponentially many combinations? We aim to optimize for an attention configuration for each transformer block under a global compute budget so as to minimize the approximated accumulated error throughout the network. More specifically, let the model have blocks indexed by , and let denote the set of candidate attention hybrid rates (e.g., ). For each block and rate , we pre-compute:
@@
This is a multiple-choice knapsack problem: select one rate per block to minimize the estimated accumulated error under a compute budget , which can be efficiently solved (Kellerer et al., 2004). The solution to this optimization identifies the final configuration of the architecture with heterogeneous attention rates that provide the best accuracy/cost trade-off under the given budget. We call the block selection homogeneous if the set consists of a single element, and heterogeneous otherwise.
@@ 4 Experimental Setup
Evaluation. We evaluate our proposed attention surgery and hybrid attention methods using the Wan2.1 1.3B video diffusion model (Wan et al., 2025). For state-of-the-art comparisons, we generate videos at the original Wan resolution and length () using the full set of extended prompts from the VBench and VBench-2.0 benchmarks (Huang et al., 2024; Zheng et al., 2025).
In addition to quantitative evaluation, we conduct a blind user study to assess visual quality and prompt alignment. We randomly select 50 prompts from VBench and present participants with paired videos, asking them to choose their preferred video or indicate no significant difference. In total, we collect 562 paired comparisons.
@@
Datasets. For fine-tuning low-resolution models, we use a 350K subset of the video dataset from Open-Sora Plan (Lin et al., 2024). For high-resolution fine-tuning, we use 22K synthetic video samples generated by Wan2.1 14B, with prompts drawn from the same source as the low-resolution dataset.
Model Hyperparameters. We experiment with converting different numbers of transformer blocks to hybrid attention: 15, 20, and 25 out of the 30 blocks in Wan2.1 1.3B. For hybrid blocks, we explore hybridization rates of 2, 4, and 8. Based on empirical analysis of the impact of and transformation complexity on generation quality, we find that a lightweight 2-layer MLP with degree-2 polynomial features is sufficient, adding approximately 2.4M parameters per converted block. Unless otherwise stated, we use value distillation loss during pretraining. Additional hyperparameter details are provided in Appendix A.
@@ 25 Blocks
| Models with 2B–5B parameters | Total | Quality | Semantic |
| Open-Sora Plan V1.3 | 77.23 | 80.14 | 65.62 |
| CogVideoX 5B | 81.91 | 83.05 | 77.33 |
| CogVideoX1.5 5B | 82.01 | 82.72 | 79.17 |
@@
| Efficient VDiT | 76.14 | – | – |
| Open-Sora V1.2 | 79.76 | 81.35 | 73.39 |
| LTX-Video | 80.00 | 82.30 | 70.79 |
| SnapGenV | 81.14 | 83.47 | 71.84 |
| Hummingbird 16frame | 81.35 | 83.73 | 71.84 |
| Wu et al. – Mobile | 81.45 | 83.12 | 74.76 |
| CogVideoX 2B | 81.55 | 82.48 | 77.81 |
| PyramidalFlow | 81.72 | 84.74 | 69.62 |
| M4V | 81.91 | 83.36 | 76.10 |
| Wan2.1 1.3B | 83.31 | 85.23 | 75.65 |
| Wan2.1 1.3B* | 82.47 | 83.33 | 79.01 |

### v3

@@ Abstract
Transformer-based video diffusion models (VDMs) deliver state-of-the-art video generation quality but are constrained by the quadratic cost of self-attention, making long sequences and high resolutions computationally expensive. While linear attention offers sub-quadratic complexity, previous approaches have failed to match the expressiveness of softmax attention unless retrained at significant computational cost. We introduce Attention Surgery, an efficient framework that enables linear or hybrid attention in pretrained VDMs, eliminating the need for training from scratch. Inspired by recent advances in language models, our method combines a novel hybrid attention mechanism—mixing softmax and linear tokens—with a lightweight distillation and fine-tuning pipeline requiring only a few GPU-days. Additionally, we incorporate a cost-aware block-rate strategy to balance expressiveness and efficiency across layers. Applied to Wan2.1 1.3B, a state-of-the-art efficient transformer VDM and evaluated on VBench, VBench2.0 and a human preference study, Attention Surgery achieves competitive results. Furthermore, measurements of on-mobile latency, memory usage, and FLOPs demonstrate notable improvements in scaling behavior for longer videos. Project page is available at: https://qualcomm-ai-research.github.io/attention-surgery
@@ 1 Introduction
Video diffusion models (VDMs) have become a cornerstone of generative modeling, enabling high-fidelity, temporally coherent video synthesis for applications from entertainment to simulation. Early VDMs relied on U-Net backbones, but these architectures struggle to scale and capture long-range temporal dependencies. Recent advances favor Diffusion Transformers (DiTs) (peebles2023dit). These models operate on spatiotemporal patches and provide global receptive fields from the outset. State-of-the-art systems such as Wan2.1 (wan2025), CogVideoX (yang2025cogvideox), HunyuanVideo (tencent2025hunyuan), PyramidalFlow (liu2025pyramidalflow), and Open-Sora Plan (lin2024opensora) exemplify this trend, consistently outperforming U-Net-based models in quality and scalability. Recent surveys confirm this transition, highlighting DiTs as the dominant architecture for video generation (wang2025survey; melnik2024survey).
While recent DiT-based video diffusion models deliver state-of-the-art quality, they come with substantial computational and memory costs that limit their practical applicability. A primary bottleneck lies in the self-attention mechanism, whose complexity scales quadratically with the sequence length, i.e., in time and in memory, where denotes the number of tokens and the hidden dimension (vaswani2017attention; rabe2021memory). This issue is particularly severe in video diffusion, where the token count easily reaches tens of thousands due to the combination of spatial patches and multiple frames. Our profiling of large-scale DiT-based video diffusion models indicates that a vast proportion of compute is devoted to self-attention. For instance, in Wan2.1 1.3B, more than 76% of the total compute within the transformer blocks is attributable to self-attention alone. Even with optimizations such as FlashAttention (dao2022flashattention), the quadratic scaling remains a fundamental barrier, constraining both training and inference when targeting higher resolutions, longer durations, or multi-shot videos. Consequently, reducing the attention cost without sacrificing model quality is critical for making video diffusion models more efficient and broadly deployable.
@@
Although linear-time attention to address the quadratic cost of softmax attention (katharopoulos2020linear; choromanski2020performer) has been around for several years, there are very few works that explore incorporating linear attention for video diffusion models. Three factors explain this gap. First, training such models from scratch is prohibitively expensive: state-of-the-art video diffusion systems require hundreds of thousands to millions of GPU hours and massive curated datasets (blattmann2023svd; chen2024videocrafter2; wan2025). Second, there is no practical method proposed for distilling softmax attention into linear attention under reasonable compute budgets for video diffusion. This difficulty arises because the exponential kernel underlying softmax requires an infinite-dimensional feature map for exact representation, making efficient approximations challenging (han2024inline; zhang2024lolcats). While linear attention in image diffusion (li2023sana) and low-rank linearization in language models (zhang2024lolcats) show promise, no analogous solution exists for the more complicated spatiotemporal token interactions in video diffusion. Third, lower expressiveness in linear attention often results in notable degradation in the attention transformation fidelity and consequently lower quality generations, specifically for videos with more complicated temporal signal dynamics.
To address these challenges, we propose an efficient attention surgery strategy – eliminating the need for extensive retraining from scratch – coupled with a novel efficient hybrid attention architecture inspired by recent developments in language modeling (zhang2024lolcats). Intuitively, if a small subset of tokens retains full softmax attention while the rest use linear attention, the model can preserve global structure and fine-grained dependencies where needed, while scaling efficiently elsewhere. Our approach significantly narrows the quality gap between linearized and full softmax attention while achieving higher efficiency than the original softmax attention models. Importantly, it can be realized with modest compute –requiring less than 0.4k GPU hours for the overall surgery – making it practical for a wide range of research and industrial settings. We validate our method on Wan2.1, a state-of-the-art video diffusion model, demonstrating that our contributions are successfully applicable to transformer-based diffusion models. Figure 2 illustrates the obtained advantage.
@@ 2 Related Work
Efficient Attention. Numerous approaches have been proposed to reduce the quadratic complexity of self-attention, for perception e.g. EfficientViT (cai2022efficientvit), PADRe (letourneaupadre), Performer (choromanski2020performer), and Linformer (wang2020linformer), for image generation e.g. SANA (li2023sana), LinGen wang2025lingen, Grafting (Grafting), and for language modeling (mercat2024linearizing; wang2024mamba; yang2024parallelizing; chen2024hedgehog; zhang2024lolcats). Although these methods demonstrate the feasibility of sub-quadratic or more efficient attention, they typically require extensive training or training from scratch (e.g., SANA (li2023sana) and SANA-Video (chen2025sana)). In contrast, our work aims for lightweight distillation and fine-tuning of pre-trained softmax-attention-based models into an efficient hybrid attention design specifically tailored for video diffusion under modest compute budgets. This direction remains valuable as the current and likely upcoming most competitive video diffusion models remain reliant on softmax attention.
@@
Linear recurrent models, such as SSM and RWKV, have recently emerged as efficient alternatives to self-attention, enabling the modeling of longer token sequences for high-resolution image generation (fei2024dimba; wang2024mamba; fei2024diffusion; zhu2025dig; yao2025diffusion). However, due to the architectural differences between transformer blocks and SSM-based blocks (e.g., Mamba), distilling pre-trained DiT weights into these architectures typically requires extensive training. In contrast, our approach preserves the same underlying block structure, enabling effective distillation under a low-training regime. Furthermore, as highlighted in the seminal work on linear attention (katharopoulos2020linear), there exists a strong connection between RNNs and causal linear attention, which allows a linearized causal attention mechanism to be deployed as an RNN during inference, a property that is particularly desirable for long video generation.
@@
Efficient Video Diffusion Models. Recent large-scale video diffusion systems such as CogVideoX (yang2025cogvideox), Open-Sora Plan (lin2024opensora), PyramidalFlow (liu2025pyramidalflow), LTX-video (hacohen2024ltx), and Wan2.1 (wan2025) have advanced quality and scalability, but at enormous compute and memory cost. Mobile-oriented designs like Mobile Video Diffusion (yahia2024mobile), MoViE (karjauv2024movie), SnapGen-V (wu2025snapgen; wu2025taming), AMD-HummingBird (isobe2025amd), and On-device Sora (kim2025device) explore lightweight architectures, yet remain non-DiT-based or rely on full quadratic attention, limiting scalability to long videos. Previous work has accelerated video generation using techniques such as token merging (bolya2023token; kahatapitiya2024object; evdit), token downsampling (crowson2024scalable; peruzzo2025adaptor), attention tiling (evdit; zhang2025fast), and sparsity (li2025compact; zhang2025faster; zhang2025spargeattn). Tiling and sparsity-based methods improve efficiency by skipping attention for most tokens. In contrast, our hybrid attention approach attends to all tokens, enabling long-range dependency modeling through a combination of linear and softmax attention. M4V (huang2025m4v) speeds up video DiTs by distilling them into Mamba blocks. We compare our model with these methods on our quantitative SOTA comparison section.
@@ 3.1 Preliminaries: Linear Attention
As observed, the terms and are independent of the output index and thus can be precomputed and cached to achieve the above-defined reformulated attention in linear complexity. Note that must be non-negative and the original linear attention paper by katharopoulos2020linear defines it as , however, the significant mismatch in expressiveness of the original similarity function and the elu-based results in substantial retraining compute and data requirement and/or inability to achieve the original quality observed from softmax attention.
@@ 3.2 Hybrid Attention
Attention Architecture. To circumvent the aforementioned issues, and inspired by the recent developments in language models (zhang2024lolcats), we define the hybrid attention by decoupling the full set of token indices into softmax tokens and the rest as linear tokens, , as:
Here is the stabilizing constant computed as the maximum exponent. As opposed to zhang2024lolcats, instead of defining as a local window around token , we uniformly subsample tokens at hybrid rate , as: . This design ensures that higher-quality softmax tokens are distributed across the entire temporal span, providing global anchors that preserve motion coherence and prevent temporal drift — an issue that local windows often suffer from in video generation. Note that selecting subsampling rates is an important hyperparameter that indicates a trade-off between the fidelity of hybrid reconstruction of the original softmax and the corresponding computational efficiency. Higher values such as or will ensure that the attention to only a small fraction of tokens scales quadratically that will notably decrease the compute burden. In contrast, a value of can more accurately reconstruct the original softmax, while still spending the quadratic terms on half of the tokens. Fig. 3 illustrates this.
@@ 3.3 Attention Surgery
where the logarithmic term mitigates numerical instabilities caused by large attention logits and gradients (barron_power_2025). However, given that matching the attention scores is a proxy optimization and the self-attention’s weighted averaged hidden states are the target to match, one can alternatively define the value distillation objective as follows:
@@
Heterogeneous Block-Rate Optimization. Referring to Fig. 4, we observe that various blocks exhibit different reconstruction error values under different hybrid rates, each with their corresponding compute costs. How do we then decide upon the hybrid rates (8, 2, 4 or 1 (full softmax)) to form our hybrid architecture among the exponentially many combinations? We aim to optimize for an attention configuration for each transformer block under a global compute budget so as to minimize the approximated accumulated error throughout the network. More specifically, let the model have blocks indexed by , and denote the set of candidate attention hybrid rates (e.g., ). For each block and rate , we pre-compute:
@@
This is a multiple-choice knapsack problem: select one rate per block to minimize the estimated accumulated error under a compute budget , which can be efficiently solved (Kellerer2004). The solution to this optimization identifies the final configuration of the architecture with heterogeneous attention rates that provide the best accuracy/cost trade-off under the given budget. We call the block selection homogeneous if the set consists of a single element, and heterogeneous otherwise.
@@ 4 Experimental Setup
Evaluation. We evaluate our proposed attention surgery and hybrid attention methods using the Wan2.1 1.3B video diffusion model (wan2025). For state-of-the-art comparisons, we generate videos at the original Wan resolution and length () using the full set of extended prompts from the VBench and VBench-2.0 benchmarks (huang2023vbench; zheng2025vbench).
In addition to quantitative evaluation, we conduct a blind user study to assess visual quality and prompt alignment. We randomly select 50 prompts from VBench and present participants with paired videos, asking them to choose their preferred video or indicate no significant difference. To prevent any random biases, the pairing order of presented videos are randomized per question. In total, we collect 562 paired comparisons.
To assess computational efficiency, we measure and report FLOPs as well as latencies and memory read/write of DiT blocks ported to Qualcomm AI Run-time (QNN) and executed on Snapdragon8-Gen4 SoC.
@@
Datasets. For fine-tuning low-resolution models, we use a 350K subset of the video dataset from Open-Sora Plan (lin2024opensora). For high-resolution fine-tuning, we use 22K synthetic video samples generated by Wan2.1 14B, with prompts drawn from the low-resolution dataset.
Model Hyperparameters. We experiment with converting different numbers of transformer blocks to hybrid attention: 15, 20, and 25 out of the 30 blocks in Wan2.1 1.3B. For hybrid blocks, we explore hybridization rates of 2, 4, and 8. Based on empirical analysis of the impact of and transformation complexity on generation quality, we find that a lightweight 2-layer MLP with degree-2 polynomial features is sufficient, adding approximately 2.4M parameters per converted block. Unless otherwise stated, we use value distillation loss during pretraining. Additional hyperparameter details are provided in the appendix.
@@ 25 Blocks
| Models with 2B–5B parameters | Total | Quality | Sem. |
| Open-Sora Plan V1.3 (lin2024opensora) | 77.23 | 80.14 | 65.62 |
| CogVideoX 5B (yang2025cogvideox) | 81.91 | 83.05 | 77.33 |
| CogVideoX1.5 5B (yang2025cogvideox) | 82.01 | 82.72 | 79.17 |
@@
| Efficient VDiT evdit | 76.14 | - | - |
| Open-Sora V1.2 zheng2024open | 79.76 | 81.35 | 73.39 |
| LTX-Video (hacohen2024ltx) | 80.00 | 82.30 | 70.79 |
| SnapGenV (wu2025snapgen) | 81.14 | 83.47 | 71.84 |
| Hummingbird (isobe2025amd) | 81.35 | 83.73 | 71.84 |
| MVDiT - Mobile (wu2025taming) | 81.45 | 83.12 | 74.76 |
| CogVideoX 2B (yang2025cogvideox) | 81.55 | 82.48 | 77.81 |
| PyramidalFlow (liu2025pyramidalflow) | 81.72 | 84.74 | 69.62 |
| M4V (huang2025m4v) | 81.91 | 83.36 | 76.10 |
| STA (zhang2025fast) | 83.00 | 85.37 | 73.52 |
| VSA (zhang2025faster) | 82.77 | 83.60 | 79.47 |
| SANA-Video (chen2025sana) | 83.71 | 84.35 | 81.35 |
| Wan2.1 1.3B (wan2025) | 83.31 | 85.23 | 75.65 |
| Wan2.1 1.3B* (wan2025) | 83.10 | 85.10 | 75.12 |